Audience profile data colected at kaap agri web site showed that 70 percent of th users were men . Assume that this percentage was based on sample 1300 users. wh is the margin of error asociated with the estimated proportion of users who are men at 95 percent confidence level



1
Expert's answer
2021-11-05T15:09:08-0400

"ME=1.96\sqrt{\frac{0.7(1-0.7)}{1300}}=0.0249."